An iron rod of length 50cm is joined to an end to an aluminium rod of length 100 cm.All measurements refer to 20°C.Find the length of the composite system at 100°C and its average,coefficient of linear expansion for iron and aluminium are 12 ×10 ^-6°C^-1 and 24×10^-6°C^-1.

Grade:11

1 Answers

Vikas TU
4 years ago
Let combined coefficient of linear expansion for the composite system be α. 
So
150(1+20α) = 50 (1+12*10^-6 *20) + 100(1+24*10^-6 * 20) 
150 + 3000α = 50+ 12*10^-3 +100+48*10^-3 
α = 20 *10^-6 /degree C 
Hope this helps 
Good Luck